# Energetic particles accelerated via turbulent magnetic reconnection in protoplanetary discs – I: ionisation rates

### v1

@@ 4.3.2 Energetic thresholds for the ionisation impact of reconnection-accelerated particles
To test the robustness of our conclusions, we vary the fraction of viscous dissipation power channelled into EP acceleration and examine the resulting ionisation rates. By reducing this fraction $\kappa$, we quantify how sensitive EP-driven ionisation is to the available energy budget and identify when EPs remain a significant source:
@@
These results show that ionisation by EPs from turbulent magnetic reconnection depends sensitively on $\kappa$. EP-dominated regions appear for efficiencies $\gtrsim 0.4\%$, and even a very small conversion of $0.025\%$ yields local contributions at the few-percent level. Moreover, although subdominant, such localized enhancements are chemically and dynamically relevant. They can modify molecular abundances, shift the ionisation balance that controls magnetic coupling, and imprint distinct line-emission features. Crucially, this behaviour differs from X-rays and GCRs, whose uncertainties (even by factors of a few) tend to affect the entire disc more uniformly. The locality of EP-driven ionisation therefore offers an observational handle. Standard GCR and X-ray rates may reproduce molecular-ion emission across much of a disc, yet still struggle to model the inner disc ionisation. Indeed, forward modelling of DM Tau (Long et al., 2024) finds that models with X-rays and GCR ionisation matching the observations in the outer disc still under-produce inner-disc HCO+, H13CO+, and N2H+ compared to observations below 50 AU, indicating an additional, more localized ionisation source in the inner regions. Accordingly, EPs accelerated by magnetic reconnection, even under conservative assumptions, should be included in thermochemical and dynamical disc models. Moreover, targeted inner-disc observations of molecular ions (e.g. HCO+, H13CO+, N2H+) can help discriminating EPs from globally acting X-rays and GCRs via their confined spatial signatures.
We also note that this study neglects (i) re-acceleration by MRI-driven turbulence (Sun and Bai, 2021) and (ii) other EP acceleration channels (Pinte et al., 2016), which could further strengthen or reshape these localized effects.
@@ 4.4 Ionisation volumetric heating rate
Ionisation not only sets chemistry but is also expected to heat the gas. When EPs ionise H or H2, energy is deposited locally through elastic/inelastic collisions and subsequent chemistry. Following the standard prescription of Glassgold et al. (2012), we take per–ionisation heat deposition $Q_{\rm H}=4.3$ eV and $Q_{\rm H_{2}}=18$ eV, and compute the EP volumetric heating rate as,
@@
This microphysically-based estimation of the heating term can be directly incorporated into thermochemical models of protoplanetary discs and outflows. Indeed, previous studies have shown that heating at the disc surface significantly affects the launching conditions of photo-evaporative and MHD-driven winds (Casse and Ferreira, 2000; Bai et al., 2016; Meskini et al., 2024).
While these studies typically include stellar X-ray and UV heating as primary sources, we now argue that they should consider EPs.
Figure 9: EP volumetric heating rate map for $\kappa=1$. The solid red contour encloses the region where energetic particles are accelerated. Dashed curves are isocontours of the EP volumetric heating rate. The upper white band marks very low columns ($N<10^{15}\mathrm{cm^{-2}}$), i.e. above the disc atmosphere.
Fig. 9 shows the spatial distribution of the EP volumetric heating rate $\Gamma_{\mathrm{EP}}$ (eV cm-3 s-1) assuming the extreme $\kappa=1$ case. Heating is the strongest in the inner disc (within a few AU) and at intermediate heights ($Z/R\sim 0.15-0.3$), where EP flux and gas density combine to maximise energy deposition, and it gradually declines toward larger radii and deeper layers. The solid red contour marks the zone where $E_{\max}>10\penalty 10000\ \mathrm{MeV}$ (efficient EP acceleration), dashed lines are isocontours, and the white region at the top indicates very low column density ($N<10^{15}\mathrm{cm^{-2}}$), above the disc atmosphere. These results indicate that EPs accelerated in the atmosphere, channel accretion and magnetic energy into heating the gas deeper in the disc.
@@
We compare the volumetric heating from EPs, $\Gamma_{\rm EP}$, to the total heating budget $\Gamma_{\rm tot}$ computed by ProDiMo (photoelectric, $\mathrm{H_{2}^{*}}$ de-excitation, $\mathrm{H_{2}}$ formation/dissociation, gas–dust accommodation, X-ray Coulomb, viscous, chemical, etc.). Fig. 10 maps the ratio $\Gamma_{\rm EP}/\Gamma_{\rm tot}$ throughout the disc. Even in the upper-limit case $\kappa=1$, EP heating is negligible near the mid-plane and generally small elsewhere (mostly $10^{-3}-10^{-2}$), becoming significant only in a narrow region at the disc surface in the inner AU and reaches close to unity values in the disc atmosphere beyond $\sim 10$ AU. For more realistic efficiencies, $\kappa\sim 0.1$, the EP contribution stays below $\sim 1\%$ across the disc, except for a small region at the surface around $R\gtrsim 10$ AU where it reaches the few percent level. These fractions are well within typical uncertainties of the total heating budget. Translated into temperature, the impact is expected to be minor compared with other heating sources. Consequently, EP heating appears secondary for thermochemical modelling, and its omission is unlikely to affect global temperature structures. By contrast, the EP ionisation term is the robust, model-relevant effect. It persists at large columns where X-rays are attenuated and can effectively modify chemistry. We therefore emphasise EP ionisation as the key addition, while treating EP heating as optional for extreme or surface-layer wind launching studies.
@@ 5.1 Turbulence-induced ionisation in discs: interpretation, and thresholds
As we have shown, EPs accelerated in situ by turbulence-induced magnetic reconnection in disc atmospheric layers can substantially reshape the ionisation structure of protoplanetary discs. When the non-thermal power is a fraction $\kappa\gtrsim 10^{-2}$, EP ionisation overtakes stellar X-rays and GCRs across extended regions, inner radii and surface/intermediate layers, while even $\kappa\sim 10^{-3}$ yields few-per-cent contributions that are still chemically relevant. Because EPs penetrate to large column densities, they maintain significant ionisation where X-rays are attenuated, providing an additional heating channel at depth.
@@
  • $\kappa\gtrsim 10^{-2}$: EPs dominated ionisation in the inner radii in the intermediate/deep layer and in the atmosphere.
  • $10^{-3}\lesssim\kappa\lesssim 10^{-2}$: EPs are subdominant yet chemically significant (few-per-cent level) in extended regions and can still matter for specific chemical tracers and non-ideal MHD terms.
  • $\kappa\lesssim 10^{-3}$: EP effects on discs can be considered insignificant.
Here $\kappa$ encapsulates the partition of released magnetic energy into heat vs. particles, and electron–proton sharing, and is likely height- and radius-dependent. Here we set $\kappa\approx 0.1$, by analogy with the canonical efficiency with which supernova-remnant shocks are thought to channel kinetic energy into GCR, in a future work, a disc-specific, first-principles estimate of $\kappa$ will be developed.
The maps are most sensitive to three ingredients: (i) the available non-thermal power $\kappa$; (ii) the ability to reach MeV energies (controlled by $\alpha_{\rm acc}$, $\tau_{\rm esc}$, $\tau_{\rm inj}$, and geometry $L_{r}/H,M_{A}$); and (iii) local gas thermodynamics through the viscous parameter $\alpha_{\rm eff}$ fixing the turbulent speed.
@@ 5.2 Implications for chemistry and observables
We expect reconnection-accelerated EPs to measurably alter the disc chemistry because their additional ionisation directly modifies the ion–molecule network and the electron fraction in the layers that they penetrate. As for X-rays and GCRs, EP ionisation of H2 initiates the $_{2}^{+}\rightarrow H+$$_{3}^{+}$ pathways, enhancing the abundance of classical ionisation tracers (e.g. HCO+, N2H+) and shifting their emitting layers. Cascades of secondary electrons further generate a CR-like internal UV field (Prasad–Tarafdar effect Prasad and Tarafdar, 1983; Padovani et al., 2024), boosting photodissociation/ionisation and grain charging at high column densities. Because the ionisation rate drives the abundance of nitriles and hydrocarbons (e.g. HCN, C2H2), introducing $\zeta_{\rm EP}(R,Z)$ where X-rays are attenuated will raise ne, push ion–molecule chemistry to larger column density, and alter abundances and lines of these tracers, yielding observable signatures. To quantify these effects, the second paper in this series of articles, will post-process our disc backgrounds with ProDiMo , adding to the standard UV/X-ray/CR sources, the EP terms $\zeta_{\rm EP}(R,Z)$ and $\Gamma_{\rm EP}(R,Z)$. We will explore a grid in non-thermal power fraction $\kappa$, spectral slope $p$, and acceleration geometry; solve to thermochemical equilibrium with/without $\zeta_{\rm EP}$; and map changes in gas temperature, electron fraction, and the layers where EPs overtake X-rays/GCRs. We will quantify EP-driven changes in classic ionisation tracers (HCO+, N2H+) and, importantly, in mid-IR prebiotic/hydrocarbon species (e.g. HCN, HNC, C2H2) that originate in the warm surface layers where EPs effect is strong, exactly the regions probed by JWST/MIRI. We will generate synthetic MIRI diagnostics (e.g. band fluxes, ratios, emitting heights, line-to-continuum ratio) to assess how EPs imprint on those spectra and how to observe them. The outcome will be abundance/temperature maps and practical discriminants that isolate EP-driven ionisation from X-ray/CR scenarios and delineate the ($\kappa,p$) domains where EP effects remain observable.
@@ 5.3 Dynamical implications for accretion and winds
First, on accretion and MRI coupling activation. Larger $n_{e}$ reduces Ohmic resistivity $\eta_{\rm O}\propto 1/n_{e}$ and, by increasing the ion density, reduces ambipolar diffusivity $\eta_{\rm A}\propto 1/\rho_{i}$. The corresponding Elsasser numbers $\Lambda_{\rm O}=v_{A}^{2}/(\eta_{\rm O}\Omega)$ and $Am=v_{A}^{2}/(\eta_{\rm A}\Omega)$ rise. Layers that were marginal (almost active, $\Lambda_{\rm O}\sim 1,Am\sim 1)$ can cross MRI thresholds, shrinking dead zones and pushing the magnetically active region to larger column densities, which will increase the local accretion rate. This creates a bootstrap process: MRI produces turbulent reconnection that accelerates EP which enhances ionisation triggering further MRI activation. In this picture, EPs produced in an MRI-active region may irradiate adjacent, marginally coupled layers, increasing the EP acceleration region size. A saturation arises from the $\beta$–dependence of EP acceleration. Turbulent-reconnection acceleration is most efficient in low-$\beta$ regions; as the MRI-active layer expands and encroaches on higher-$\beta$ gas, the reconnection-driven EP production drops. Once the front of the MRI-active zone reaches these high-$\beta$ layers, it can no longer sustain a supra-thermal EP population able to ionise, so the outward growth of the acceleration region naturally stalls, quenching the bootstrap loop. A related pathway may operate in regions where the field is still sufficiently frozen-in to wind up. Reversals between toroidal field loops can generate thin current sheets even when MRI modes are suppressed. These sheets can undergo bursts of turbulent reconnection, producing rapid EP acceleration and temporary spikes in ionisation, potentially activating MRI and EP acceleration locally and intermittently.
@@
Second, on wind launching and magneto-thermal coupling. Magneto-thermal winds require gas that is well coupled to a reasonably coherent large-scale field at the launch base. Disc layers where winds are anchored are expected to be turbulent and only partially ionised, which lowers conductivities and suppress efficient wind launching solutions. By adding $\zeta_{\rm EP}(R,Z)$ where X-rays fade, EPs can thicken the well-coupled layer and, through their volumetric heating $\Gamma_{\rm EP}$ aiding mass loading. This does not guarantee large-scale field coherence, but it may broaden the parameter space in which a steady magneto-thermal wind can be maintained, shifting the launch region and altering mass-loading and torque.
@@
In a future work, we will post-process this additional EP source on a ProDiMo disc structure to clarify the importance and spatial extent of this bootstrap process and whether this mechanism can patch or reduce the dead zone. However, only dedicated non-ideal MHD simulations, including turbulent reconnection and EP transport self-consistently, can properly assess the impact of this mechanism. Including $\zeta_{\rm EP}(R,Z)$ and $\Gamma_{\rm EP}(R,Z)$ in conductivity and energy modules will modify vertical profiles of $\eta_{\rm O}$, $\eta_{\rm A}$ (and Hall where relevant), the MRI-active layer thickness, and the wind-launch region. Because EPs penetrate to larger column density than X-rays, their impact is concentrated exactly where simulations currently struggle to maintain adequate coupling between the disc material and magnetic field. While EPs will not by themselves “solve” the wind-launching problem, they provide an additional, physically motivated ionisation and heating channel that can tip marginal layers into the coupled regime, with measurable consequences for both accretion and outflows.

### v2

@@ 4.3.2 Energetic thresholds for the ionisation impact of reconnection-accelerated particles
To test the robustness of our conclusions, we varied the fraction of viscous dissipation power channelled into EP acceleration and examined the resulting ionisation rates. By reducing this fraction $\kappa$, we quantified how sensitive EP-driven ionisation is to the available energy budget and identified when EPs remain a significant source:
@@
These results show that ionisation by EPs from turbulent magnetic reconnection depends sensitively on $\kappa$. EP-dominated regions appear for efficiencies $\gtrsim 0.4\%$, and even a very small conversion of $0.025\%$ yields local contributions at the level of a few percent. Moreover, although subdominant, such localised enhancements are chemically and dynamically relevant. They can modify molecular abundances, shift the ionisation balance that controls magnetic coupling, and imprint distinct line-emission features. Crucially, this behaviour differs from X-rays and GCRs, whose uncertainties (even by factors of a few) tend to affect the entire disc more uniformly. The locality of EP-driven ionisation therefore offers an observational handle. Standard GCR and X-ray rates may reproduce molecular-ion emission across much of a disc, yet still struggle to model the inner disc ionisation. Indeed, forward modelling of DM Tau (Long et al., 2024) finds that models with X-rays and GCR ionisation matching the observations in the outer disc still under-produce inner-disc HCO+, H13CO+, and N2H+ compared to observations below 50 AU, indicating an additional, more localised ionisation source in the inner regions. Accordingly, EPs accelerated by magnetic reconnection, even under conservative assumptions, should be included in thermochemical and dynamical disc models. Moreover, targeted inner-disc observations of molecular ions (e.g. HCO+, H13CO+, N2H+) can help to discriminate EPs from globally acting X-rays and GCRs via their confined spatial signatures. We also note that this study neglects (i) re-acceleration by MRI-driven turbulence (Sun and Bai, 2021) and (ii) other EP acceleration channels (Pinte et al., 2016), which could further strengthen or reshape these localised effects.
@@ 4.4 Ionisation volumetric heating rate
Ionisation not only sets chemistry but is also expected to heat the gas. When EPs ionise H or H2, energy is deposited locally through elastic and inelastic collisions and subsequent chemistry. Following the standard prescription of Glassgold et al. (2012), we took per–ionisation heat deposition $Q_{\rm H}=4.3$ eV and $Q_{\rm H_{2}}=18$ eV, and computed the EP volumetric heating rate as
@@
This microphysically-based estimation of the heating term can be directly incorporated into thermochemical models of protoplanetary discs and outflows. Indeed, previous studies have shown that heating at the disc surface significantly affects the launching conditions of photo-evaporative and MHD-driven winds (Casse and Ferreira, 2000; Bai et al., 2016; Meskini et al., 2024). While these studies typically include stellar X-ray and UV heating as primary sources, we now argue that they should consider EPs.
Figure 9: EP volumetric heating rate map for $\kappa=1$. The solid red contour encloses the region where EPs are accelerated. Dashed curves are isocontours of the EP volumetric heating rate. The upper white band marks very low columns ($N<10^{15}\mathrm{cm^{-2}}$), i.e. above the disc atmosphere.
Fig. 9 shows the spatial distribution of the EP volumetric heating rate, $\Gamma_{\mathrm{EP}}$ (eV cm-3 s-1), assuming the extreme $\kappa=1$ case. Heating is the strongest in the inner disc (within a few astronomical units) and at intermediate heights ($Z/R\sim 0.15-0.3$), where EP flux and gas density combine to maximise energy deposition, and it gradually declines toward larger radii and deeper layers. The solid red contour marks the zone where $E_{\max}>10\penalty 10000\ \mathrm{MeV}$ (efficient EP acceleration), dashed lines are isocontours, and the white region at the top indicates a very low column density ($N<10^{15}\mathrm{cm^{-2}}$), above the disc atmosphere. These results indicate that EPs accelerated in the atmosphere channel accretion and magnetic energy into heating the gas deeper in the disc.
@@
We compared the volumetric heating from EPs, $\Gamma_{\rm EP}$, to the total heating budget, $\Gamma_{\rm tot}$, computed by ProDiMo (photoelectric, $\mathrm{H_{2}^{*}}$ de-excitation, $\mathrm{H_{2}}$ formation and dissociation, gas–dust accommodation, X-ray Coulomb, viscous, chemical, etc.). Fig. 10 maps the ratio $\Gamma_{\rm EP}/\Gamma_{\rm tot}$ throughout the disc. Even in the upper-limit case $\kappa=1$, EP heating is negligible near the mid-plane and generally small elsewhere (mostly $10^{-3}-10^{-2}$), becoming significant only in a narrow region at the disc surface in the inner AU and reaches close to unity values in the disc atmosphere beyond $\sim 10$ AU. For more realistic efficiencies, $\kappa\sim 0.1$, the EP contribution stays below $\sim 1\%$ across the disc, except for a small region at the surface around $R\gtrsim 10$ AU where it reaches the few percent level. These fractions are well within typical uncertainties of the total heating budget. Translated into temperature, the impact is expected to be minor compared with other heating sources. Consequently, EP heating appears secondary for thermochemical modelling, and its omission is unlikely to affect global temperature structures. By contrast, the EP ionisation term is the robust, model-relevant effect. It persists at large columns where X-rays are attenuated and can effectively modify chemistry. We therefore emphasise EP ionisation as the key addition, while treating EP heating as optional for extreme or surface-layer wind launching studies.
@@ 5.1 Turbulence-induced ionisation in discs: Interpretation, and thresholds
As we have shown, EPs accelerated in situ by turbulence-induced magnetic reconnection in disc atmospheric layers can substantially reshape the ionisation structure of protoplanetary discs. When the non-thermal power is a fraction $\kappa\gtrsim 10^{-2}$, EP ionisation overtakes stellar X-rays and GCRs across extended regions, inner radii and surface to intermediate layers, while even $\kappa\sim 10^{-3}$ yields contributions of a few percent that are still chemically relevant. Because EPs penetrate to large column densities, they maintain significant ionisation where X-rays are attenuated, providing an additional heating channel at depth.
@@
  • $\kappa\gtrsim 10^{-2}$: The EPs dominated ionisation in the inner radii in the intermediate/deep layer and in the atmosphere.
  • $10^{-3}\lesssim\kappa\lesssim 10^{-2}$: The EPs are subdominant yet chemically significant (at the level of a few percent) in extended regions and can still matter for specific chemical tracers and non-ideal MHD terms.
  • $\kappa\lesssim 10^{-3}$: The EP effects on discs can be considered insignificant.
Here $\kappa$ encapsulates the partition of released magnetic energy into heat versus particles, and electron–proton sharing, and is likely height- and radius-dependent. Here we set $\kappa\approx 0.1$, by analogy with the canonical efficiency with which supernova-remnant shocks are thought to channel kinetic energy into GCR, in a future work, a disc-specific, first-principles estimate of $\kappa$ will be developed.
The maps are most sensitive to three ingredients: (i) the available non-thermal power, $\kappa$; (ii) the ability to reach mega-electronvolt energies (controlled by $\alpha_{\rm acc}$, $\tau_{\rm esc}$, $\tau_{\rm inj}$, and geometry $L_{r}/H,M_{A}$); and (iii) local gas thermodynamics through the viscous parameter, $\alpha_{\rm eff}$, fixing the turbulent speed.
@@ 5.2 Implications for chemistry and observables
We expect reconnection-accelerated EPs to measurably alter the disc chemistry because their additional ionisation directly modifies the ion–molecule network and the electron fraction in the layers that they penetrate. As for X-rays and GCRs, EP ionisation of H2 initiates the $_{2}^{+}\rightarrow H+$$_{3}^{+}$ pathways, enhancing the abundance of classical ionisation tracers (e.g. HCO+, N2H+) and shifting their emitting layers. Cascades of secondary electrons further generate a CR-like internal UV field (Prasad–Tarafdar effect Prasad and Tarafdar, 1983; Padovani et al., 2024), boosting photodissociation, ionisation and grain charging at high column densities. Because the ionisation rate drives the abundance of nitriles and hydrocarbons (e.g. HCN, C2H2), introducing $\zeta_{\rm EP}(R,Z)$ where X-rays are attenuated will raise ne, push ion–molecule chemistry to larger column density, and alter abundances and lines of these tracers, yielding observable signatures. To quantify these effects, the second paper in this series of articles, will post-process our disc backgrounds with ProDiMo , adding to the standard UV/X-ray/CR sources, the EP terms $\zeta_{\rm EP}(R,Z)$ and $\Gamma_{\rm EP}(R,Z)$. We will explore a grid in non-thermal power fraction $\kappa$, spectral slope $p$, and acceleration geometry; solve to thermochemical equilibrium with and without $\zeta_{\rm EP}$; and map changes in gas temperature, electron fraction, and the layers where EPs overtake X-rays/GCRs. We will quantify EP-driven changes in classic ionisation tracers (HCO+, N2H+) and, importantly, in mid-IR prebiotic and hydrocarbon species (e.g. HCN, HNC, C2H2) that originate in the warm surface layers where EPs effect is strong, exactly the regions probed by JWST/MIRI. We will generate synthetic MIRI diagnostics (e.g. band fluxes, ratios, emitting heights, line-to-continuum ratio) to assess how EPs imprint on those spectra and how to observe them. The outcome will be abundance and temperature maps and practical discriminants that isolate EP-driven ionisation from X-ray/CR scenarios and delineate the ($\kappa,p$) domains where EP effects remain observable.
@@ 5.3 Dynamical implications for accretion and winds
First, on accretion and MRI coupling activation, larger $n_{e}$ reduces Ohmic resistivity $\eta_{\rm O}\propto 1/n_{e}$ and, by increasing the ion density, reduces ambipolar diffusivity, $\eta_{\rm A}\propto 1/\rho_{i}$. The corresponding Elsasser numbers, $\Lambda_{\rm O}=v_{A}^{2}/(\eta_{\rm O}\Omega)$ and $Am=v_{A}^{2}/(\eta_{\rm A}\Omega)$, rise. Layers that were marginal (almost active, $\Lambda_{\rm O}\sim 1,Am\sim 1)$ can cross MRI thresholds, shrinking dead zones and pushing the magnetically active region to larger column densities, which will increase the local accretion rate. This creates a bootstrap process: MRI produces turbulent reconnection that accelerates EP, which enhances ionisation triggering further MRI activation. In this picture, EPs produced in an MRI-active region may irradiate adjacent, marginally coupled layers, increasing the EP acceleration region size. A saturation arises from the $\beta$–dependence of EP acceleration. Turbulent-reconnection acceleration is most efficient in low-$\beta$ regions; as the MRI-active layer expands and encroaches on higher-$\beta$ gas, the reconnection-driven EP production drops. Once the front of the MRI-active zone reaches these high-$\beta$ layers, it can no longer sustain a supra-thermal EP population able to ionise, so the outward growth of the acceleration region naturally stalls, quenching the bootstrap loop. A related pathway may operate in regions where the field is still sufficiently frozen-in to wind up. Reversals between toroidal field loops can generate thin current sheets even when MRI modes are suppressed. These sheets can undergo bursts of turbulent reconnection, producing rapid EP acceleration and temporary spikes in ionisation, potentially activating MRI and EP acceleration locally and intermittently.
@@
Second, on wind launching and magneto-thermal coupling, magneto-thermal winds require gas that is well coupled to a reasonably coherent large-scale field at the launch base. Disc layers where winds are anchored are expected to be turbulent and only partially ionised, which lowers conductivities and suppresses efficient wind launching solutions. By adding $\zeta_{\rm EP}(R,Z)$ where X-rays fade, EPs can thicken the well-coupled layer and, through their volumetric heating, $\Gamma_{\rm EP}$, aid mass loading. This does not guarantee large-scale field coherence, but it may broaden the parameter space in which a steady magneto-thermal wind can be maintained, shifting the launch region and altering mass-loading and torque.
@@
In a future work, we will post-process this additional EP source on a ProDiMo disc structure to clarify the importance and spatial extent of this bootstrap process and whether this mechanism can patch or reduce the dead zone. However, only dedicated non-ideal MHD simulations, including turbulent reconnection and EP transport self-consistently, can properly assess the impact of this mechanism. Including $\zeta_{\rm EP}(R,Z)$ and $\Gamma_{\rm EP}(R,Z)$ in conductivity and energy modules will modify vertical profiles of $\eta_{\rm O}$, $\eta_{\rm A}$ (and Hall where relevant), the MRI-active layer thickness, and the wind-launch region. Because EPs penetrate to a larger column density than X-rays, their impact is concentrated exactly where simulations currently struggle to maintain adequate coupling between the disc material and magnetic field. While EPs will not by themselves ’solve’ the wind-launching problem, they provide an additional, physically motivated ionisation and heating channel that can tip marginal layers into the coupled regime, with measurable consequences for both accretion and outflows.